# Exciton dissociation in organic solar cells: An embedded charge transfer state model

### v1

@@ Abstract
Organic solar cells are a promising avenue for renewable energy, and our study introduces a comprehensive model to investigate exciton dissociation processes at the donor-acceptor interface. Examining quantum efficiency and emitted phonons in the charge transfer state (CTS), we explore scenarios like variations of the environment beyond the CTS and repulsive/attractive potentials. The donor-acceptor interface significantly influences the injection process, with minimal impact from the environment beyond the CTS. Attractive potentials can create localized electron states at the interface, below the acceptor band, without necessarily hampering a good injection at higher energies. Exploring different recombination processes, including acceptor-side and donor-side recombination, presents distinct phases in the energy-injection versus recombination rate. Our study highlights the important role of the type of recombination in determining the quantum efficiency and the existence of hot or cold charge transfer states. Finally, depending on the initial energy of the electron on the donor side, three distinct injection regimes are exhibited. The present model should be helpful for optimizing organic photovoltaic cell interfaces, highlighting the critical parameter interplay for enhanced performance.
@@ 1 Introduction
A significant limiting factor toward the development of more efficient OSCs is the lack of a comprehensive understanding of their working principles. For instance, despite significant efforts on both the experimental and theoretical sides, the mechanism responsible for exciton dissociating into free charges is still poorly understood. Various phenomena have been recognized as aiding in charge separation (i.e., exciton dissociation), including built-in electric fields at donor-acceptor (D:A) interfaces, the delocalization of excitons and charge carriers, energetic variability, and structural disorder [10, 7, 8, 9, 11, 12, 13]. Beyond these crucial aspects related to electronic states at the interface and electrostatic effects, the charge transfer process is dynamic. Consequently, it is important to describe the dynamics of this process, such as the excitation of vibrational modes or the competition between charge separation and charge recombination between the electron and the hole at the charge transfer state (CTS). The CTS is the initial state in the acceptor where the electron hops and becomes coupled to local vibrational modes.
@@
On the theoretical side, various numerical methods have been proposed for a fully microscopic understanding of the charge separation mechanism, including Monte Carlo simulations [14, 15], exact diagonalization [16], and time-dependent density functional theory [17, 18, 19]. Additionally, several models have been developed, such as the Braun-Onsager analytic model [20, 21], the Marcus theory [22] and its related approaches, and ab initio electronic structure calculations [23, 24, 25]. However, none of these models have adequately clarified the processes occurring at the donor-acceptor interface during charge transfer. New complementary approaches are needed.
@@
In this article, we propose a quantum model for charge separation at a donor-acceptor interface, focusing on electron-phonon coupling and the partially coherent nature of the electron. The central concept is the CTS, where the electron moves to the acceptor side and couples with local vibration modes. This coupling allows the electron to excite phonons, leading to a vibrationally hot CTS if multiple phonons are excited, or a cold CTS if none are. The efficiency of charge separation is debated in relation to energy release on these modes [11, 12, 13, 26, 27, 28, 29, 30, 31]. In this embedded CTS model, we take also into account other effects such as environmental factors and electrostatic potential. We treat the dynamics of the transfer exactly by combining the quantum scattering theory and the dynamical Mean Field Theory (DMFT) [16, 32, 33, 34, 35]. The recombination is also treated on a quantum basis. This allows for the treatment of the strong coupling regime between the electron and vibrational modes, in contrast to perturbative theories.
@@
The paper is organized as follows. We begin with a presentation of the embedded CTS model and then discuss briefly the formalism that allows to compute the output of the dynamical injection process. The physics of the injection process is analyzed firstly for simple cases and then with the full complexity of the model. We then proceed to a global discussion of the results where we distinguish three regimes according to the value of the initial energy of the electron.
@@ 2.3 Recombination and electrostatic potential
There are several electron-recombination processes with the hole on the donor site. The electron, when it is on the donor site in the excited state, can recombine with the hole in the ground state, also located on the donor site. Similarly, when it injects itself onto the acceptor site, the electron can recombine with the hole it left behind. Figure 3 illustrates the two recombination processes: on the left, recombination from the donor site, and on the right, recombination from the acceptor site.
The recombination mechanism can involve a complex sequence of quantum processes, but as we consider the zero-temperature limit, we treat recombination as the injection of the electron into a band. If recombination is accompanied by the emission of a photon, the continuum of states accessible by this recombination is broad in energy compared to a typical polaronic bandwidth. However, other recombination processes may occur for which the continuum of states may not necessarily be broad compared to a typical polaronic bandwidth. Thus, we distinguish between different types of injection corresponding to these possibilities, and we show that this can have importance on the course of injection.
@@
There are two types of potential at the donor/acceptor interface: either an attractive potential between the electron on the CTS site and the hole on the donor site, or a potential (potentially repulsive) between the electron on the CTS and the negative charges of the molecules in the donor.
@@ 2.4 Holstein Hamiltonian
The total Hamiltonian $H$ takes the form [36, 37]:
@@
Where the index $i$ represents distinct sites, with $i=0$ denoting the CTS. $\varepsilon_{I}$ stands for the energy of the incoming electron of a molecule located at the donor site. The hopping parameter $m$ between the donor site and the CTS is considered weak compared to the pure electronic bandwidth $4J$, allowing us to approach the limit $m\to$ 0. For a given site $i$, the creation (annihilation) operators for electrons are denoted as $c_{i}^{+}$ ($c_{i}$). The electrostatic potential at site $i$, situated on layer $L_{i}$, is expressed as $\frac{V}{L_{i}+1}$, resulting from electron interaction with the hole or with interfacial charges and characterized by the parameter $V$. The creation (annihilation) operators for the local vibration mode at site $i$ are represented by $a_{i}^{+}$ ($a_{i}$), with $\hbar\omega$ representing the phonon energy ( we consider $\hbar=1$) . The hopping matrix elements $J_{i,j}$ between nearest neighbors $i$ and $j$ on the Bethe lattice are expressed from $J$. In our approach, we adopt the standard limit of infinite coordination $K$ for the Bethe lattice, while simultaneously imposing a finite bandwidth of $4J$. As $K$ tends to infinity, the term $K(J_{i,j})^{2}$ stabilizes to a constant value, converging towards $J^{2}$. This implies that each hopping integral $J_{i,j}$ diminishes as $K$ increases, ensuring that the bandwidth remains constant. $g$ signifies the electron-vibration coupling parameter (we used the same phonon mode for all sites, $g_{i}=g$). The Hamiltonian $H_{R}$ describes the recombination processes. We work in the limit where the temperature is zero, because band energies and phonon energies are well above the thermal energy at room temperature. All energies are expressed in units of $J$.
@@ 3.1 Channels and probabilities
From a $|n\rangle$ state, there are two channels through which the electron leaves the CTS. The $nA$ channel corresponds to propagation on the acceptor side, and the $nR$ channel corresponds to the recombination process. The probabilities of injection into the $nA$ ($nR$) channels are $\Phi_{A}^{n}$ ($\Phi_{R}^{n}$). The probabilities $\Phi_{A}^{n}$ and $\Phi_{R}^{n}$ of injection into the $nA$ and $nR$ channels are calculated from scattering theory and depend only on the self-energies $\Delta_{A}(\varepsilon_{I}-n\hbar\omega)$ and $\Delta_{R}(\varepsilon_{I}-n\hbar\omega)$ that describe the corresponding channels. The quantity $\Delta_{A}(\varepsilon_{I}-n\hbar\omega)$ corresponds to the injection self-energy that depends on the model for the acceptor. In contrast, $\Delta_{R}(\varepsilon_{I}-n\hbar\omega)$ represents the recombination self-energy. To calculate the self-energies, we use the recursion method coupled to the DMFT as shown in the reference [38].
@@ 5 Numerical study
In this part, we analyze complex situations with all the possible effects included in the embedded model such as electron-vibration coupling, electrostatic potential, the role of the environment, role of the type of recombination process. We start by analyzing the electronic structure through the density of states (DOS) on the CTS $|0>$. This part is independent of the recombination which has a negligible influence on the DOS. Then we analyze the yield and energy transfer on the CTS. We consider first a few cases for some representative parameter sets. We continue with a presentation of the phase diagram as a function of two essential parameters which are the injection energy and the recombination rate. These phase diagrams confirm the importance of the nature of the dominant recombination process for the yield and the energy transfer on the CTS.
@@ 5.1 Electronic structure: role of environment and electrostatic potential
In the case of an attractive potential $V=-2$, Figure 9 c) shows that the density of states presents localized peaks at the band bottom. Model B is again more sensitive, although the general trends are consistent. Note that the total weight of the localized on the state $|0>$ can be quite high of the order of $0.3-0.4$ for $g=1$.
Now, let’s focus on the ground state of the polaron in the presence of an attractive potential, which corresponds to a localized polaronic state.
@@ 5.2 Quantum Yield and energy transfer on the CTS
Figure 11 presents the quantum yield and the average phonon energy for parameters $g=1$, $\omega=0.8$, and $\Gamma_{R}=0.1$. The figure includes three panels: with no potential, with a repulsive potential ($V=2$), and with an attractive potential ($V=-2$). Model A is shown in red, and model B is shown in green. The recombination is treated in this case as injection in a wide band.
@@ 6 Discussion
Within the framework of the embedded CTS model and for the range of parameters considered here, we ultimately distinguish three main regimes for injection based on the value of the injection energy $\varepsilon_{I}$. In order of increasing $\varepsilon_{I}$, we find a no-injection regime, a resonant injection regime, and a tunneling injection regime. We summarize all the results in Figure 17.
@@
In a higher energy range, the injection is called resonant. $\varepsilon_{I}$ lies within the continuum of states, allowing quick injection from the donor to the CTS and then rapid electron transfer to the acceptor without phonon excitation on the CTS (cold CTS). We found that the electrostatic potential in this regime does not qualitatively alter the behavior. Notably, even if an attractive potential can create a bound state with a significant weight on the CTS (> 0.3-0.4), it does not prevent efficient injection into the continuum of states. Thus, the existence of bound states created by the electron-hole interaction in the acceptor is not incompatible with good injection and high quantum yield. In this resonant regime, the effect of the environment beyond the CTS is also moderate.
In an even higher energy range, we find a tunneling regime. For the electron to be injected into the environment beyond the CTS, its initial energy $\varepsilon_{I}$ must first be reduced by exciting phonons on the CTS to resonate with the acceptor band. This situation is analogous to a potential barrier, and the concept of Wigner time for barrier crossing (the time required for energy transfer on the CTS) helps analyze the behavior of the electron + CTS system. In this regime injection into the acceptor is accompanied by a hot CTS since the electron must lose energy before injection in the acceptor band. Additionally, as the density of states on the CTS decreases, the injection time from the donor increases, according to Fermi’s golden rule. Thus, in this tunneling regime, the residence times on the donor and on the CTS are long, and if recombination processes are fast enough, they will reduce the quantum yield of the injection and possibly lead to a cold CTS. Note that in the model of narrow-band recombination, the recombination time also increases with $\varepsilon_{I}$. Therefore, once injected into the CTS, the electron may still have a significant injection yield on the acceptor side (see Figure 11 b)).
@@ 7 Conclusion
In this article, we have developed and studied a charge injection model at a donor-acceptor interface, typically applicable in organic semiconductors. This model, which we refer to as the "embedded charge transfer state model", allows for the analysis of the injection process by focusing on the electron transfer to the CTS site and simplifying the environment on both the acceptor and donor sides. We have employed a framework based on dynamic mean-field theory and scattering theory, accounting for electron-lattice interactions, electrostatic potential on the acceptor side, and geminate recombination between electron and hole.
The study emphasizes that the efficiency of charge injection in organic solar cells is strongly influenced by the interplay between injection and recombination processes. It shows that the coupling between electrons and vibrational modes affects quantum efficiency, introducing the concepts of hot and cold charge transfer states. Our findings indicate that interface parameters are crucial for optimizing cell performance and that the interface’s impact is more significant than environmental variations beyond the CTS. In addition, the electrostatic potential can play a moderate role in the injection process, even when it is attractive and creates localized bound electron-hole states at the interface.
Our analysis highlights three primary injection regimes based on the electron injection energy $\varepsilon_{I}$: a no-injection regime at low $\varepsilon_{I}$, an efficient resonant injection regime at intermediate $\varepsilon_{I}$ with a cold CTS, and a tunneling injection regime at higher $\varepsilon_{I}$ with a more difficult injection into the acceptor which is accompanied by a hot CTS.
We believe that the present model offers a simplified but comprehensive approach. It should be helpful for a better understanding of the exciton dissociation process and for finding ways to improve the organic solar cells.

### v2

@@ Abstract
Organic solar cells are a promising avenue for renewable energy, and our study introduces a comprehensive model to investigate exciton dissociation processes at the donor-acceptor interface. Examining quantum efficiency and emitted phonons in the charge transfer state (CTS), we explore scenarios like variations of the environment beyond the CTS and repulsive/attractive potentials. The donor-acceptor interface significantly influences the injection process, with minimal impact from the environment beyond the CTS. Attractive potentials can create localized electron states at the interface, below the acceptor band, without necessarily hampering a good injection at higher energies. Exploring different recombination processes, including acceptor-side and donor-side recombination, presents distinct phases for the injection process versus the initial energy of the electron and the recombination rate. Our study highlights the important role of the type of recombination in determining the quantum efficiency and the existence of hot or cold charge transfer states. Finally, depending on the initial energy of the electron on the donor side, three distinct injection regimes are exhibited. The present model should be helpful for optimizing organic photovoltaic cell interfaces, highlighting the critical parameter interplay for enhanced performance.
@@ 1 Introduction
A significant limiting factor toward the development of more efficient OSCs is the lack of a comprehensive understanding of their working principles. For instance, despite significant efforts on both the experimental and theoretical sides, the mechanism responsible for exciton dissociating into free charges is still poorly understood. Various phenomena have been recognized as aiding in charge separation (i.e., exciton dissociation), including built-in electric fields at donor-acceptor (D:A) interfaces, the delocalization of excitons and charge carriers, energetic variability, and structural disorder [10, 7, 8, 9, 11, 12, 13]. Beyond these crucial aspects related to electronic states at the interface and electrostatic effects, the charge transfer process is dynamic. Consequently, it is important to describe the dynamics of this process, such as the excitation of vibrational modes or the competition between charge separation and charge recombination between the electron and the hole [14].
@@
On the theoretical side, various numerical methods have been proposed for a fully microscopic understanding of the charge separation mechanism, including Monte Carlo simulations [15, 16], exact diagonalization [17], and time-dependent density functional theory [18, 19, 20]. Additionally, several models have been developed [11, 21], such as the Braun-Onsager analytic model [22, 23], the Marcus theory [24] and its related approaches, and ab initio electronic structure calculations [25, 26, 27]. However, none of these models have adequately clarified the processes occurring at the donor-acceptor interface during charge transfer. New complementary approaches are needed.
@@
In this article, we propose a quantum model for charge separation at a donor-acceptor interface, focusing on electron-phonon coupling and the partially coherent nature of the electron. The central concept is the CTS, where the electron moves to the acceptor side and couples with local vibration modes. This coupling allows the electron to excite phonons, leading to a vibrationally hot CTS if multiple phonons are excited, or a cold CTS if none are. The efficiency of charge separation is debated in relation to energy release on these modes [11, 12, 13, 28, 29, 30, 31, 32, 33].A first analysis of charge injection was presented in our previous work [34]. Here we enlarge our approach by considering several types of recombination processes, various type of environments, attractive or repulsive potential and by analysing more deeply the process of injection and its interdependence with phonon emission on the CTS. In the context of the model our numerical treatment of the dynamics combines the quantum scattering theory and the dynamical Mean Field Theory (DMFT) and is exact.[17, 35, 36, 37, 38].This allows for the treatment of the strong coupling regime between the electron and vibrational modes, in contrast to perturbative theories.
@@
The paper is organized as follows. We begin with a presentation of the embedded CTS model and then discuss briefly the formalism that allows to compute the output of the dynamical injection process. The physics of the injection process is analyzed firstly for simple cases and then with the full complexity of the model. We then proceed to a global discussion of the results where we distinguish three regimes for injection, according to the value of the initial energy of the electron.
@@ 2.3 Recombination and electrostatic potential
There are several electron-recombination processes with the hole on the donor site. The electron, when it is on the donor site in the excited state, can recombine with the hole in the ground state, also located on the donor site. Similarly, when it injects into the acceptor site, the electron can recombine with the hole it left behind. Figure 3 illustrates the two recombination processes: on the left, recombination from the donor site, and on the right, recombination from the acceptor site.
Several mechanisms can lead to the genimate recombination between the electron in the acceptor side and the hole left in the donor side. This can be for example photon emission or multiple emission of low energy phonons which lead to a transfer of the electron into localized states near the interface and then to recombination with the hole [14]. If recombination is accompanied by the emission of a photon, the continuum of states accessible by this recombination is broad in energy compared to a typical polaronic bandwidth. However, for the other recombination processes the continuum of excitations may not necessarily be broad compared to a typical polaronic bandwidth. Thus, we distinguish between these two different types of injection, and we show that this can have importance on the course of injection.
@@
There are two types of potential at the donor/acceptor interface: either an attractive potential V<0 between the electron on the CTS site and the hole on the donor site, or a repulsive potential V>0 between the electron on the CTS and the negative charges of the molecules in the donor.
@@ 2.4 Holstein Hamiltonian
The Hamiltonian $H$ of our model takes into account a tight-binding hopping model for the electron and includes the essential coupling with local vibration modes, as is common for molecular solids [21]. $H$ is of a Holstein type [39, 40] and takes the form :
@@
Where the index $i$ represents distinct sites, with $i=0$ denoting the CTS. $\varepsilon_{I}$ stands for the energy of the incoming electron of a molecule located at the donor site. The hopping parameter $m$ between the donor site and the CTS is considered weak compared to the pure electronic bandwidth $4J$, allowing us to approach the limit $m\to$ 0. For a given site $i$, the creation (annihilation) operators for electrons are denoted as $c_{i}^{+}$ ($c_{i}$). The electrostatic potential at site $i$, situated on layer $L_{i}$, is expressed as $\frac{V}{L_{i}+1}$, resulting from electron interaction with the hole or with interfacial charges and characterized by the parameter $V$. The creation (annihilation) operators for the local vibration mode at site $i$ are represented by $a_{i}^{+}$ ($a_{i}$), with $\hbar\omega$ representing the phonon energy ( we consider $\hbar=1$) . The hopping matrix elements $J_{i,j}$ between nearest neighbors $i$ and $j$ on the Bethe lattice are expressed from $J$. In our approach, we adopt the standard limit of infinite coordination $K$ for the Bethe lattice, while simultaneously imposing a finite bandwidth of $4J$. As $K$ tends to infinity, the term $K(J_{i,j})^{2}$ stabilizes to a constant value, converging towards $J^{2}$. This implies that each hopping integral $J_{i,j}$ diminishes as $K$ increases, ensuring that the bandwidth remains constant. $g$ signifies the electron-vibration coupling parameter (we used the same phonon mode for all sites, $g_{i}=g$). The Hamiltonian $H_{R}$ describes the recombination processes. As stated above these processes can be represented by a coupling to a continuum of states and in the formalism used here we just need to model the self-energy $\Delta_{R}(z)$ that represents this coupling to a continuum. We consider the cases where the continuum is a wide band ($\Delta_{R}(z)=-i\Gamma_{R}$) and narrow band as discussed in Section 5.2. All energies are expressed in units of $J$, which is typically of the order of $0.2$ eV. All parameters are chosen to fall within a realistic experimental range [41, 32, 42, 31, 30, 43, 44]. Finally we work in the limit where the temperature is zero, because band energies and phonon energies are well above the thermal energy at room temperature.
@@ 3.1 Channels and probabilities
From a $|n\rangle$ state, there are two channels through which the electron leaves the CTS. The $nA$ channel corresponds to propagation on the acceptor side, and the $nR$ channel corresponds to the recombination process. The probabilities of injection into the $nA$ ($nR$) channels are $\Phi_{A}^{n}$ ($\Phi_{R}^{n}$). The probabilities $\Phi_{A}^{n}$ and $\Phi_{R}^{n}$ of injection into the $nA$ and $nR$ channels are calculated from scattering theory and depend only on the self-energies $\Delta_{A}(\varepsilon_{I}-n\hbar\omega)$ and $\Delta_{R}(\varepsilon_{I}-n\hbar\omega)$ that describe the corresponding channels. The quantity $\Delta_{A}(\varepsilon_{I}-n\hbar\omega)$ corresponds to the injection self-energy that depends on the model for the acceptor. In contrast, $\Delta_{R}(\varepsilon_{I}-n\hbar\omega)$ represents the recombination self-energy. To calculate the self-energies $\Delta_{A}(z)$ due to the coupling with the acceptor, we use the recursion method coupled to the DMFT as shown in the reference [34].
@@ 5 Numerical study
In this part, we analyze complex situations with all the possible effects included in the embedded model such as electron-vibration coupling, electrostatic potential, the role of the environment, role of the type of recombination process. We start by analyzing the electronic structure through the density of states (DOS) on the CTS $|0>$. This part is independent of the recombination which has a negligible influence on the DOS. Then we analyze the yield and energy transfer on the CTS. We consider first a few cases for some representative parameter sets. We continue with a presentation of the phase diagram as a function of two essential parameters which are the injection energy and the recombination rate. These phase diagrams confirm the importance of the nature of the dominant recombination process for the yield and the energy transfer on the CTS. Note that the parameters of the Hamiltonian are given in units of $J$ which is typically of the order of $0.2$ eV. All parameters are chosen such that they are in a realistic experimental range to model the prototypical PCMB and C60 acceptor systems [41, 32, 42, 31, 30, 43, 44].These molecular solids have electronic bands which are narrow (typically 4J which is less than one eV) and have molecular vibration modes of high frequency which can be of the order of 0.1/0.2 eV.
@@ 5.1 Electronic structure: role of environment and electrostatic potential
In the case of an attractive potential $V=-2$, Figure 9 c) shows that the density of states presents localized peaks at the band bottom. Model B is again more sensitive, although the general trends are consistent. Note that the total weight of the localized on the state $|0>$ can be quite high of the order of $0.3-0.4$ for $g=1$. These states are analogous to bound electronic states in atoms or impurity states in semiconductors and are geometrically localized around the CTS. Although an infinite series of such states might be expected near the minimum energy of the continuum, their weights are too low to be detected numerically.
Now, let’s focus on the ground state of the polaron in the presence of an attractive potential, which corresponds to a localized polaronic state (Model A).
@@ 5.2 Quantum Yield and energy transfer on the CTS
Figure 11 presents the quantum yield and the average phonon energy for parameters $g=1$, $\omega=0.8$, and $\Gamma_{R}=0.1$. The figure includes three panels: with no potential, with a repulsive potential ($V=2$), and with an attractive potential ($V=-2$). Model A is shown in red, and model B is shown in green. The recombination is treated in this case as injection in a wide band ($\Delta_{R}=-i\Gamma_{R}$).
@@ 6 Discussion
Within the framework of the embedded CTS model and for the range of parameters considered here, we ultimately distinguish three main regimes based on the value of the injection energy $\varepsilon_{I}$. In order of increasing $\varepsilon_{I}$, we find a no-injection regime, a resonant injection regime, and a tunneling injection regime. We summarize all the results in Figure 17.
@@
In a higher energy range, the injection is resonant. Indeed $\varepsilon_{I}$ lies within the continuum of states, allowing quick injection from the donor to the CTS and then rapid electron transfer to the acceptor without phonon excitation on the CTS (cold CTS). We found above that a repulsive potential seems more favorable to a good injection with a cold CTS than an attractive potential, even though the electrostatic potential does not qualitatively alter the injection behavior. Notably, even if an attractive potential can create a bound state with a significant weight on the CTS (> 0.3-0.4), it does not prevent efficient injection into the continuum of states. Thus, the existence of bound states created by the electron-hole interaction in the acceptor is not incompatible with good injection and high quantum yield. In this resonant regime, the effect of the environment beyond the CTS is also moderate.
In an even higher energy range, we find a tunneling regime. For the electron to be injected into the environment beyond the CTS, its initial energy $\varepsilon_{I}$ must first be reduced by exciting phonons on the CTS to get in resonance with the acceptor band. This situation is analogous to the crossing of a potential barrier, and the concept of Wigner time for barrier crossing can be applied here. It indicates that in this regime the electron spends more time on the CTS and can therefore be more sensitive to recombination. Additionally, as the density of states on the CTS decreases at high energies $\varepsilon_{I}$, the injection time from the donor increases, according to Fermi’s golden rule. Thus, in this tunneling regime, the residence times of the electron on the donor side and on the CTS are longer. Therefore if recombination processes are fast enough, they will reduce the quantum yield of the injection and possibly lead to a cold CTS. In the wide band recombination model the time for recombination is constant and will always become smaller than the Wigner time for residence on the CTS when $\varepsilon_{I}$ increases. Yet in the narrow-band recombination model, the recombination time also increases with $\varepsilon_{I}$. Therefore, once injected into the CTS, the electron may still have a significant injection yield on the acceptor side (see Figure 11 b)).
@@ 7 Conclusion
In this article, we have developed and studied a charge injection model at a donor-acceptor interface, typically applicable in organic semiconductors. This model, which we refer to as the "embedded charge transfer state model", allows for the analysis of the injection process by focusing on the electron transfer to the CTS site and simplifying the environment on both the acceptor and donor sides. This model accounts for electron-lattice interactions, electrostatic potential on the acceptor side, and geminate recombination between electron and hole. We have employed a framework based on dynamic mean-field theory and scattering theory, which provides us with an exact numerical solution of this model.
This study shows that the coupling between electrons and vibrational modes affects quantum efficiency, introducing the concepts of hot and cold charge transfer states. Our findings indicate that interface parameters are crucial for optimizing cell performance and that the interface’s impact is more significant than environmental variations beyond the CTS. In addition, the electrostatic potential can play a moderate role in the injection process, even when it is attractive and creates localized bound electron-hole states at the interface. The charge injection process is described as a transfer into a continuum at an energy $\varepsilon_{I}$. When $\varepsilon_{I}$, is in the bottom part of the continuum an efficient resonant injection regime occurs with a cold CTS. At higher values of $\varepsilon_{I}$ the electron must first loose energy by emitting phonons before going in the acceptor side. In this tunneling regime the electron spends more time on the donor orbital and then on the CTS site and the recombination can be favored.
Although a detailed comparison with a given system is beyond the scope of this work, we believe that the present model offers a simplified but comprehensive approach to advanced analyses of experimental results [11]. It should be helpful for a better understanding of the exciton dissociation process and for finding ways to improve the organic solar cells. The present embedded charge transfer state model could still be improved by considering several orbitals and several modes per site and even finite temperature. We emphasize that the modeling of the dynamics of the recombination process is also an essential ingredient.